The long-term liability section of Johnson Laboratories balance sheet as of December 31, 2014, included 10% bonds having a face amount of $230 million and a remaining premium of $9.8 million. On January 2, 2015, Johnson retired some of the bonds before their scheduled maturity. Required: Prepare the journal entry by Johnson to record the redemption of the bonds under each of the independent circumstances below:

1. Johnson called half the bonds at the call price of 103.6.

2. Johnson repurchased $92 million of the bonds on the open market at their market price of $97.5 million.
Answer #1

Given values are

10% Bonds Payable           = $230 Million

Premium on bonds payable = $9.8 Million

Working note: Interest payable for 2 days= ($230*10%)*2/365=0.13 $ Million. As the Interest is very minimal and it will get set off with Premium balance in the future. So, The above interest (calculated) amount is ignored in the below solutions.

  1. Called half the Bonds at the price of 103.6

Then, Total amount payable for above redemption (including Premium)

= ($230/2)*103.6/100=$119.14 Millions

Premium on redemption = 119.14-115=$4.14 Million

*Redemption entry for the above call is:-

            Bonds Payable   a/c                           -Dr      $115 Million

            Premium on Bonds payable a/c      -Dr       $4.14 Million

                                      To    Cash/Bank a/c                 $119.14 Million.

(Being half bonds redeemed at a premium has been accounted).

  1. Repurchase of $92 Million of the Bonds at $97.5 million

In the above offer, Premium paying on redemption (repurchase) = 97.5-92= $5.5 Million

*Entry for this redemption/repurchase in the books of account as follows:-

             Bonds Payable   a/c                           -Dr      $92 Million

            Premium on Bonds payable a/c      -Dr       $5.5 Million

                                      To    Cash/Bank   a/c                 $97.5 Million.

(Being Bonds redeemed at the Market price has been accounted).